OpenStudy (anonymous):

solve. 4/7 = x-3 over 21

OpenStudy (anonymous):

try to get all x to one side and all numbers to other side

OpenStudy (anonymous):

\[\frac{4}{7} = \frac{x-3}{21}\] \[\frac{4}{1} = \frac{x-3}{3}\] \[{4}\times{3} = {x-3}\] \[{12} = {x-3}\] \[{12+3} = {x-3+3}\] \[{15} = {x}\]